Fred is 3 times old than his sister. Four years ago he was 5 times as old as his sister. Find their present age.
Answer by macston(5194)

.
S=sister's age; F=Fred's age=3S
.
F-4=5(S-4) Substitute for F.
3S-4=5S-20
16=2S
8=S ANSWER 1: Sister is 8 years old,
F=3S=3(8)=24 ANSWER 2: Fred is 24 years old.
.
CHECK:
.
F-4=5(S-4)
24-4=5(8-4)
20=5(4)
20=20